OpenStudy (anonymous):

2(-q-7)-5=13 in steps

OpenStudy (lgbasallote):

try transposing 5 :)

OpenStudy (anonymous):

what

OpenStudy (anonymous):

what does tranposing mean?

OpenStudy (lgbasallote):

uhmm i meant putting the 5 on the other side by adding 5 to both sides

OpenStudy (anonymous):

2(-q-7)=8

OpenStudy (anonymous):

-q-7= 9 -q=16 q=-16?

OpenStudy (lgbasallote):

that's 2(-q-7) = 13 + 5

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

then

OpenStudy (lgbasallote):

@Romero transpose means "To reverse or transfer the order or place of" so i meant putting 5 on other side and changing the side :D

OpenStudy (lgbasallote):

sign*

OpenStudy (anonymous):

what do i do now

OpenStudy (lgbasallote):

you have 2(-q - 7) = 18 divide both sides by 2

OpenStudy (anonymous):

(-q-7)=9

OpenStudy (lgbasallote):

now add 7 to both sides

OpenStudy (anonymous):

-q=16

OpenStudy (lgbasallote):

now divide both sides by -1 to make your q positive

OpenStudy (anonymous):

q=-16

OpenStudy (lgbasallote):

there you go

OpenStudy (anonymous):

how old are you cause you are really smart

OpenStudy (anonymous):

7(p+3)-4p=-9 srry about that question it is rude to ask a lady how old she is

OpenStudy (lgbasallote):

haha lol it's fine :)) distribute 7 first

OpenStudy (anonymous):

what i am very bad at the vocab

OpenStudy (lgbasallote):

multiply 7 to (p+3) each

OpenStudy (anonymous):

7p+21-4p=-9

OpenStudy (anonymous):

hello lgb

OpenStudy (lgbasallote):

alright..now combine 7p and - 4p

OpenStudy (anonymous):

11p+21=-9

OpenStudy (anonymous):

lgb

OpenStudy (lgbasallote):

subtract 21 from both sides

OpenStudy (anonymous):

11p=-30

OpenStudy (lgbasallote):

now divide both sides by 11

OpenStudy (anonymous):

p=-2.7

OpenStudy (lgbasallote):

right!